OpenStudy (anonymous):

Simple question Series by @rohangrr A ship is moving at a speed of 56 km/h. one sec later, it is moving at 58 km/hr. what is its acceleration.??

OpenStudy (anonymous):

@ujjwal @lgbasallote (ambassador) @PeX

OpenStudy (anonymous):

welcome

ganeshie8 (ganeshie8):

120 km/h^2

OpenStudy (shane_b):

If it went from 56 km/hr to 58km/h then its acceleration is simply 2km/h/s

OpenStudy (anonymous):

Point To Remember: 1.Velocity and time are related as \[\quad v=u+at \] 2.position and time are related as \[s=ut+\frac { 1 }{ 2 } at^{ 2 }\quad \] 3.position and velocity are related as \[{ v }^{ 2 }-{ u }^{ 2 }=2as\]

OpenStudy (ujjwal):

\[\frac{58-56}{\frac{1}{60\times 60}} km/h^2\]

OpenStudy (anonymous):

okay, then

OpenStudy (anonymous):

explain properly

OpenStudy (ujjwal):

Acceleration= change in velocity/time

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

then

OpenStudy (anonymous):

every body felt asleep!1!!!

OpenStudy (ujjwal):

LOL..

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

@Vaidehi09 i asked u smthing right??

OpenStudy (ujjwal):

I gave the formula.. I have put the values earlier. what is not clear to you?

OpenStudy (anonymous):

its clear lol

OpenStudy (anonymous):

I knew the answer but still for sake of asking u all lolll

OpenStudy (anonymous):

vaidehi alive

OpenStudy (anonymous):

@Rohangrr yes...replying now: well I'll be a first year B.Tech student starting this august. and answer to the abv ques: is it 5/9 m/s^2 ?

OpenStudy (anonymous):

yes @ujjwal whos right

OpenStudy (anonymous):

if u know the answer then why r u posting it?

OpenStudy (anonymous):

well u know on my dash board theres a tag kept ASK U R FIRST QUESTION

OpenStudy (anonymous):

:\ then ask a worthwhile ques

OpenStudy (anonymous):

then you will be not able to answer it

OpenStudy (ujjwal):

Both answers are right. @Rohangrr

OpenStudy (anonymous):

lol thats why i gave u medal lol

OpenStudy (anonymous):

i never knew there was a competition going on here.

OpenStudy (anonymous):

yes !!! highly competitive

OpenStudy (ujjwal):

what competition?

OpenStudy (anonymous):

dude that medal and levels

OpenStudy (anonymous):

those are not for competing with others. u fail to see the whole point of this system.

OpenStudy (ujjwal):

LOL... That doesn't define you! Even the one who has spent more time has higher level. A higher lever doesn't mean you are genius!

OpenStudy (anonymous):

\[\Huge\mathcal{W}\mathcal{a}\mathcal{i}\mathcal{t}\]

OpenStudy (anonymous):

i mean vaidahi all izzz well a half empty glass is always half full is how u see

OpenStudy (anonymous):

top discussing level bck to physics